OpenStudy (anonymous):

A sphere has a surface area of 36 pi ft^2. Find the volume of the sphere

zepdrix (zepdrix):

Surface Area of a sphere is given by, \[\Large A=4\pi r^2\] They told us our Surface Area is 36pi.\[\Large 36\pi=4\pi r^2\] Solve for \(\Large r\). Once you've successfully done that, plug this new found radius into the Volume Formula,\[\Large V=\frac{4}{3}\pi r^3\]

OpenStudy (anonymous):

How would I solve for r?

zepdrix (zepdrix):

\[\Large 36\pi=4\pi r^2\]Dividing both sides by \(\large \pi\) gives us,\[\Large 36=4r^2\]Dividing each side by 4 gives us,\[\Large \frac{36}{4}=r^2\]Simply, then take the square root of both sides to solve for r :o

OpenStudy (anonymous):



OpenStudy (anonymous):



zepdrix (zepdrix):

Yay good job \c:/